OpenStudy (anonymous):

A wave length has a frequency of 6.1 x 10^14 Hz and a speed of 3.00 x 10^8 m/s. What is the wavelength?

OpenStudy (anonymous):

This is in the wrong section (should be in physics). However, I can assist. Wavelength Equation: λ=v/f Where: v=Phase speed f=Frequency Solve: 3.00E8/6.1E14=4.91E21

OpenStudy (anonymous):

Thank you, I didn't know there was a specific subject to put the question in.

OpenStudy (anonymous):

It's ok haha

OpenStudy (anonymous):

@Starak_47 Hai Sweet Friend, WeLcoMe to OpeN stuDy.. !! Wondering If you are satisfied with this answer, Please close this Question. Thank You ! Keep in touch with @LoVe_ByTeS Bye !